OpenStudy (anonymous):

The expression log a^5b is equivalent to... a. 5 log a + log b b. 5 (log a + log b) c. log 5 a + log b d. log 5 ab

Parth (parthkohli):

\[\log(a^{5b})= 5b \log a\]

OpenStudy (aravindg):

@danadacnomania first apply property \[\log a + \log b=\log ab \]